$SPAD/src/input bug101.input

Timothy Daly December 3, 2016

Abstract

Contents

2

The call

laplace(log(z),z,w)

causes an infinite loop between lfintegrate and monomialIntegrate.

Waldek modified the code to fail and return unevaluated.

A more reasonable return value, not supported by Axiom, would be:

  -log(w)-gamma
  -------------
       w

— * —

)set break resume
)spool bug101.output
)set message test on
)set message auto off
)clear all
--S 1 of 2
laplace(log(z),z,w)
--R
--R   (1)  laplace(log(z),z,w)
--R                                                    Type: Expression(Integer)
--E 1
--S 2 of 2
laplace(log(z),w,z)
--R
--R        log(z)
--R   (2)  ------
--R           z
--R                                                    Type: Expression(Integer)
--E 2

———-

— * —

)spool
)lisp (bye)

———-

References^{[1] nothing}

4